17. Find the period of the following function: y = sin (2x) cos (6x) + cos (2x) sin (6x)
Answer #1

y=sin(2x) cos(6x) + cos(az)sin (62) = sin(2x+62) [:sinAcosB + costsinB = sin(A+8)] y =sin (8x) period of sina is at period of